Urban Window Company had gross wages of $320,000 during the week ended July 15. The amount of wages subject to social security tax was $320,000, while the amount of wages subject to federal and state unemployment taxes was $40,000. Tax rates are as follows:
Social security 6.0%
Medicare 1.5%
State unemployment 5.4%
Federal unemployment 0.6%
The total amount withheld from employee wages for federal taxes was $75,200.
Required:
A. Journalize the entry to record the payroll for the week of July 15.*
B. Journalize the entry to record the payroll tax expense incurred for the week of July 15.*
*Refer to the Chart of Accounts for exact wording of account titles.

A. Journalize the entry to record the payroll for the week of July 15.*
B. Journalize the entry to record the payroll tax expense incurred for the week of July 15.*
*Refer to the Chart of Accounts for exact wording of account titles.
Answer #1
Date Description Post Ref. Debit Credit Assets Liabilities Equity
July 15 Wages expense 320000
       Social Security tax payable 19200 =320000*6%
       Medicare tax payable 4800 =320000*1.5%
       Employees federal Income tax payable 75200
       Wages payable 220800
July 15 Payroll tax expense 26400
       Social Security tax payable 19200
       Medicare tax payable 4800
       Federal Unemployment tax payable 240 =40000*0.6%
       State Unemployment tax payable 2160 =40000*5.4%